yoselinr2006:

Two friends wash cars to make extra money. The profit P(x) of one friend after x days can be represented by the function P(x) = −x2 + 9x + 10. The second friend's profit can be determined by the function Q(x) = 12x. Solve the system of equations. What solution is a viable answer to the question, "After how many days will the two students earn the same profit?" and which solution is a nonviable answer? Show your work and justify your answer.

yoselinr2006:

i really need help on itttt

jhonyy9:

can you solve a system of equations ?

yoselinr2006:

I'm not really good at it i don't fully understand it.. so I'm having trouble trying to answer this one

yoselinr2006:

I do flvs..and my algebra2 grade is a d... and i need to do a discussion assignment on the question..and I need help

jhonyy9:

\[-x^{2} + 9x +10 = 0\] \[12x = 0\] this is your system

yoselinr2006:

okk

yoselinr2006:

how do i explain the strategy tho

yoselinr2006:

because i have to explain how i gotthat answer...and the answer i put before was from the intertnet.. i used the wrong equation

jhonyy9:

ok I will try explain to you - hope understandably 12x = 0 how you calcule the x value from this equation ?

yoselinr2006:

you find a number to substitute 0 and then divide them ??

jhonyy9:

divide both sides by 12 so in this way what value will get for x ?

yoselinr2006:

you would be left with just x

yoselinr2006:

so 0?

jhonyy9:

yes of cours x = 0

yoselinr2006:

okk so thats rigt

jhonyy9:

ok remember on this please so now the first equation is a quadratic function - yes ?

yoselinr2006:

yes

jhonyy9:

ok what is the diagram of a quadratic function ?

jhonyy9:

is a parabola

jhonyy9:

two equations when has common roots ?

yoselinr2006:

its a quadratic function..

jhonyy9:

when their diagrams intersects and the intersection point is the common root when two equations havent common roots so from this result that their diagram dont intersect never

yoselinr2006:

ohh

yoselinr2006:

okk

jhonyy9:

in this case for second equation you get x = 0 and hope you know that this equation has a diagram what intersect the origin yes ?

yoselinr2006:

yes

jhonyy9:

so now then you substitute the x=0 in place of x inside the first equation so what will get ? the left side value will be equal with right side value of 0 ?

jhonyy9:

please try it

yoselinr2006:

12?..im sorryy

jhonyy9:

no for P(0) = ? what will get ?

yoselinr2006:

and if not than i think 1

jhonyy9:

@jhonyy9 wrote:
\[-x^{2} + 9x +10 = 0\] \[12x = 0\] this is your system
here in place of x substitute 0 p(0) = ?

yoselinr2006:

i think its something divided by 12 a numbe greatrer than 1 ofc..maybee 6

yoselinr2006:

im not sure

jhonyy9:

\[P(0)= -0^{2} +9×0 +10\] so P(0) = ?

yoselinr2006:

19 to the power 0f 2 ?can that give me the answer

yoselinr2006:

if 9 times 0 is 0 then its just 9 + 10

yoselinr2006:

9+10 - 0^2

jhonyy9:

no the first term there is - zero on power of two = ?

yoselinr2006:

0

jhonyy9:

@jhonyy9 wrote:
\[P(0)= -0^{2} +9×0 +10\] so P(0) = ?
here the first term -0 on power of 2 = 0 yes ? the second term how you said is correct so 9×0 = 0 and now there are 0 + 0 +10 = ?

yoselinr2006:

jus 10

jhonyy9:

from this result P(0) = 10 and you remember the second equation was Q(0) = 0 yes ?

yoselinr2006:

right

jhonyy9:

so for the same x = 0 these equations give different results from this result that these equations havent common roots so what mean that their diagram dont intersect never what result from this like answer on your posted problem ?

yoselinr2006:

wha do you mean

jhonyy9:

@xxdeppressionxx

yoselinr2006:

so the answer is x=0

yoselinr2006:

and the equations do not intetrsect

jhonyy9:

no the answer not is x= 0 bc. this is a root just for second equation but from this result that these equations havent common roots so dont will intersect never and what mean this for these two friends ? can they have the same profit ?

yoselinr2006:

no

yoselinr2006:

because they have no common root

jhonyy9:

because these two equations havent common roots so this mean that these two friends not can have the same profit never

jhonyy9:

Idk do you understand my help ?

yoselinr2006:

yes, just a Lil...but thank you so much for your time I appreciate it.

jhonyy9:

ok np anytime cya

WickedHeart:

@yoselinr2006 If this is done please close